OpenStudy (anonymous):

Given the function f(x) = −2x + 4, find x if f(x) = 20. −8 −36 44 −12 show steps for solving

OpenStudy (anonymous):

-2x+4=20 -2x=-24 x=12

OpenStudy (anonymous):

isnt it -12

OpenStudy (caozeyuan):

f(x) =20 so -2x+4=20 so -2x=16 so x=-8, not 12!!!! @chandanjha

OpenStudy (anonymous):

so its -8 im so confused

OpenStudy (caozeyuan):

20-4=16!

OpenStudy (caozeyuan):

16/-2=-8!

OpenStudy (caozeyuan):

sorry, there is no factorial

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

yeah i was wrong thnks

OpenStudy (caozeyuan):

yeah, A

OpenStudy (anonymous):

yeah a

OpenStudy (anonymous):

thanks for the verification